# Gradient descent induces alignment between weights and the empirical NTK for deep non-linear networks

### v1

@@ 4 Increasing the centered contribution to the uncentered NFA improves feature learning
Our theoretical and experimental work has established that gradient based training leads to alignment of the weight matrices to the PTK feature covariance. This process is driven by the C-NFA. Therefore, one path towards improving the correlations defined by the NFA is to increase the contribution of the C-NFA to the dynamics - as measured, for example, by the ratio $\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)\text{tr}\left(W^{\top}WW^{\top}KW\right)^{-1}$. When this ratio is large, the C-NFA contributes significantly in magnitude to the UC-NFA, indicating successful feature learning. We will discuss two interventions which increase this ratio, with an eye towards improving generalization.
@@ 4.1 Feature learning and initialization
(a) Standard training.
(b) Speed Limited Optimization.
Figure 4: The uncentered and centered NFA for the first layer weight matrix as a function of initialization, with and without SLO. We consider a two hidden layer network with ReLU activations, where we set $C_{0}=500$, and $C_{1}=C_{2}=0.002$. The third column shows the ratio of the unnormalized C-NFA  correlation to the UC-NFA: ${\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)\cdot{\text{tr}\left(W^{\top}WW^{\top}KW\right)}^{-1}}$. We show that SLO pushes the ratio of the unnormalized C-NFA to UC-NFA close to $1$ and improves the UC-NFA across initialization scales.
One factor that modulates the level of feature learning is the scale of the initialization in each layer. In our experiments training networks with unmodified gradient descent, we observe that the centered NFA will increasingly dominate the uncentered NFA with decreasing initialization (third panel, first row, Figure 4). Further, as the centered NFA increases in contribution to the uncentered quantity, the strength of the UC-NFA, and to a lesser extent, the C-NFA increases (Figure 4). The decreases in correspondence between these quantities is also associated with a decrease in the feature quality of the NFM (Appendix H) and test error (Figure 2) for the chain monomial task.
For fully-connected networks with homogeneous activation functions, such as ReLU, and no normalization layers, decreasing initialization scale is equivalent to decreasing the scale of the outputs, since we can write $f(Wx)=a^{-p}f(aWx)$ for any scalar $a$ for any homogenous activation $f$. This in turn is equivalent to increasing the scale of the labels. Therefore, decreasing initialization forces the weights to change more in order to fit the labels, leading to more change in $F$ from its initialization. Conceptually, this aligns with the substantial line of empirical and theoretical evidence that increasing initialization scale or output scaling transitions training between the lazy and feature learning regimes (Chizat et al., 2019; Woodworth et al., 2020; Agarwala et al., 2020; Lyu et al., 2023).
This relationship suggests that small initialization can be broadly applied to increase the change in F, and the strength of the UC-NFA. However, this may not be ideal; for example, if the activation function is differentiable at 00, small initialization leads to a network which is approximately linear. This may lead to low expressivity unless the learning dynamics can increase the weight magnitude.
@@ 4.2 Speed Limited Optimization
(a) Standard training.
(b) SLO.
Figure 5: The NFM and AGOP from a trained network with and without Speed Limited Optimization with fixed initialization scale on the chain monomial task. With standard training and large initialization, the NFM remains close to identity, and does not learn task structure. SLO increases feature learning, causing the NFM to become highly correlated with the AGOP and learn the target features.
We can instead design an intervention which can increase feature learning without the need to decrease the initialization scale. We do so by fixing the learning speed layerwise to constant values, which causes the C-NFA to dominate the UC-NFA dynamics. For weights at layer $\ell$ and learning rate $\eta>0$, we introduce Speed Limited Optimization (SLO), which is characterized by the following update rule,
@@
where the hyperparameter $C_{\ell}\geq 0$ controls the amount of learning in layer $\ell$. We expect this rule to increase the strength of the UC-NFA in layers where $C_{\ell}$ is large relative to $C_{m}$ for $m\neq\ell$, as $W^{(\ell)}$ will be forced to change significantly from initialization. As a result, $\left\|W_{0}^{(\ell)}\right\|\left\|W_{t}^{(\ell)}\right\|^{-1}\rightarrow 0$ for large $t$, causing the centered and uncentered NFA to coincide for this layer.
We demonstrate the effects of SLO on the chain-monomial task (Equation (1)). We found that fixing the learning speed the be high in the first layer and low in the remaining layers causes the ratio of the unnormalized C-NFA to the UC-NFA to become close to $1$ across initialization scales (Figure 4b). The same result holds for the SVHN dataset (see Appendix K). We note that this intervention can be applied to target underperforming layers and improve generalization in deeper networks (see Appendix J for details).
We observe that the correlations from both the C-NFA and the UC-NFA become close to $1$ after training with SLO, independent of initialization scale (Appendix H). Further, the quality of the features learned, measured by the similarity of $F$ and $\bar{G}$ to the true EGOP, significantly improve and are more similar to each other with SLO (Figure 5), even with large initialization. In contrast, in standard training the UC-NFA fails to develop with large initialization, as $F$ resembles identity (no feature learning), while $\bar{G}$ only slightly captures the relevant features.
@@ 5 Discussion
Centered NFA through training
Our theoretical analysis in this work shows that the PTK feature covariance at initialization has a relatively simple structure in terms of the weights and the data. To predict the correlation of the NFA later in training, we will likely need to account for the change in this matrix. One should be able to predict this development at short times by taking advantage of the fact that eigenvectors of the Hessian change slowly during training (Bao et al., 2023). An alternative approach would be to use a quadratic model for neural network dynamics (Agarwala et al., 2022; Zhu et al., 2022) which can capture dynamics of the NTK (and therefore the PTK).
@@
We provide an analytic prediction for the C-NFA at initialization for one hidden layer networks with quadratic activation functions, and provide first-order predictions for two hidden layer networks with quadratic and ReLU activations. A followup direction would be to use random matrix analysis to more accurately predict the NFA for deep architectures with general activations. A promising approach in this direction is to approximate non-linear activations with a linear expansion and independent correction terms (Adlam et al., 2019; Adlam and Pennington, 2020; El Karoui, 2008; Hu and Lu, 2022).
More generally it’s an open question as to how the NFA develops with depth in more complicated architectures like CNNs and transformers. More research is needed to understand how PTK-weight alignment relates to generalization and trainability in deep networks.
@@
There are prior works that implement differential learning rates across layers (Howard and Ruder, 2018; Singh et al., 2015), though these differ from differential learning speeds as in SLO, in which the L2-norm of the gradients are fixed, and not the scaling of the gradients. Our intervention is more similar to the LARS (You et al., 2017) and LAMB (You et al., 2019) optimizers that fix learning speeds to decrease training time in ResNet and BERT architectures. However, these optimizers fix learning speeds to the norm of the weight matrices, while our intervention sets learning speeds to free hyperparameters in order to increase feature learning, as measured by the NFA. We also note as an additional explanation for the success of SLO in our experiments, the PTK feature covariance may change slowly (or just in a single direction) with this optimizer, because the learning rate is small for later layers, allowing the first layer weights to align with the PTK.
@@
We demonstrate in this work that the dataset (by our construction in Section 3) and optimizer (with SLO in Section 4) play a significant role in the strength of feature learning. Important future work would be to understand other settings where significant empirical differences exist between optimization choices. In particular, analyzing the NFA may clarify the role of gradient batch size and adaptive gradient methods in generalization (Zhu et al., 2023).
6 Acknowledgements
We thank Lechao Xiao for detailed feedback on the manuscript. We also thank Jeffrey Pennington for helpful discussions.
Appendix A Omitted proofs of propositions in the main text
@@
Proof of Proposition 6.
At initialization, we have, $\dot{W}=\frac{\partial f(X)}{\partial h_{1}}^{\top}\dot{\mathcal{L}}X$. Therefore, using that $\mathcal{K}\equiv\frac{\partial f(X)}{\partial h_{1}}\frac{\partial f(X)}{\partial h_{1}}^{\top}$,
$\dot{W}^{\top}\dot{W}=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h_{1}}\frac{\partial f(X)}{\partial h_{1}}^{\top}\dot{\mathcal{L}}X=X^{\top}\dot{\mathcal{L}}\mathcal{K}\dot{\mathcal{L}}X~{},$
and,
$\dot{W}^{\top}K\dot{W}$ $=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h_{1}}\,K\,\frac{\partial f(X)}{\partial h_{1}}^{\top}\dot{\mathcal{L}}X$
$=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h_{1}}\frac{\partial f(X)}{\partial h_{1}}^{\top}\frac{\partial f(X)}{\partial h_{1}}\frac{\partial f(X)}{\partial h_{1}}^{\top}\dot{\mathcal{L}}X$
$=X^{\top}\dot{\mathcal{L}}\mathcal{K}^{2}\dot{\mathcal{L}}X~{}.$
Appendix B Additional centerings of the NFA
Double-centered NFA
One may additionally center the PTK feature map to understand the co-evolution of the PTK feature covariance and the weight matrices. In this work, we consider such a centering that we refer to as the double-centered NFA.
Ansatz 7 (Double-centered NFA).
$(\bar{W}^{(\ell)})^{\top}\bar{W}^{(\ell)}\propto(\bar{W}^{(\ell)})^{\top}\bar{K}^{(\ell)}\bar{W}^{(\ell)}~{},$
where $\bar{K}^{(\ell)}=\left(\frac{\partial f(X)}{\partial h_{\ell}}-\frac{\partial f_{0}(X)}{\partial h_{\ell}}\right)^{\top}\left(\frac{\partial f(X)}{\partial h_{\ell}}-\frac{\partial f_{0}(X)}{\partial h_{\ell}}\right)$, and $f_{0}$ is the neural network at initialization.
However, the double-centered NFA term corresponds to higher-order dynamics that do not significantly contribute the centered and uncentered NFA (Figure 6) when initialization is large or for early periods of training. Note however this term becomes relevant over longer periods of training.
Figure 6: Ratio of the unnormalized double-centered NFA correlation to the centered NFA correlation throughout neural network training. In particular, we plot $\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}\bar{K}\bar{W}\right)\cdot{\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)}^{-1}$ throughout training for both layers of a two-hidden layer MLP with ReLU activations.
@@ Appendix B Additional centerings of the NFA
One may also center just the PTK feature map, while substituting the initial weights for $W$ to isolate how the PTK feature covariance aligns to the weight matrices. To measure this alignment, we consider the PTK-centered NFA.
Ansatz 8 (PTK-centered NFA).
$(W_{0}^{(\ell)})^{\top}W_{0}^{(\ell)}\propto(W_{0}^{(\ell)})^{\top}\bar{K}^{(\ell)}W_{0}^{(\ell)}~{},$
where $W_{0}^{(\ell)}$ is the initial weight matrix at layer $\ell$.
@@
Figure 7: PTK-centered NFA correlation throughout training for both layers of a two-hidden layer MLP with ReLU activations on Gaussian data with two different spectra.
Appendix C Extending our theoretical predictions to depth and general activations
Precise predictions of the C-NFA become more complicated with additional depth and general activation functions. However, we note that the deep C-NFA will remain sensitive to a first-order approximation in which $K$ is replaced by its expectation. We demonstrate that this term qualitatively captures the behavior of the C-NFA for 2 hidden layer architectures with quadratic and, to a lesser extent, ReLU activation functions in Figure 8. In this experiment, we sample Gaussian data with mean 00 and covariance with a random eigenbasis. We parameterize the eigenvalue decay of the covariance matrix by a parameter $\alpha$, called the data decay rate, so that the eigenvalues have values $\lambda_{k}=\frac{1}{1+k^{\alpha}}$. As $\alpha$ approaches 00 or $\infty$ the data covariance approaches a projector matrix.
In this experiment, we see that the data covariance spectrum will also parameterize the eigenvalue decay of $\mathbb{E}\left[K\right]$, allowing us to vary how close the expected PTK matrix (and its dual, the PTK feature covariance) is from a projector, where the NFA holds exactly. We see that for intermediate values of $\alpha$, both the observed and the predicted derivatives of the C-NFA decreases in value.

### v2

@@ 5 Increasing the centered contribution to the NFC strengthens feature learning
Our theoretical and experimental work has established that gradient based training leads to alignment of the weight matrices to the PTK feature covariance. This process is driven by the C-NFC. Therefore, one path towards improving the neural feature correlations is to increase the contribution of the C-NFC to the dynamics - as measured, for example, by the centered-to-uncentered, or, C/UC, ratio $\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)\text{tr}\left(W^{\top}WW^{\top}KW\right)^{-1}$. When this ratio is large, the C-NFC contributes significantly in magnitude to the UC-NFC, indicating successful feature learning as measured by the uncentered NFC. We will discuss how small initialization promotes feature learning, and design an optimization rule, Speed Limited Optimization, that increases the C/UC ratio and drives the value of the UC-NFC to $1$.
@@ 5.1 Feature learning and initialization
One factor that modulates the level of feature learning is the scale of the initialization in each layer. In our experiments training networks with unmodified gradient descent, we observe that the centered NFC will increasingly dominate the uncentered NFC with decreasing initialization (third panel, first row, Figure 4). Further, as the centered NFC increases in contribution to the uncentered quantity, the strength of the UC-NFC, and to a lesser extent, the C-NFC increases (Figure 4). The decreases in correspondence between these quantities is also associated with a decrease in the feature quality of the NFM (Appendix H), for the chain monomial task.
For fully-connected networks with homogeneous activation functions, such as ReLU, and no normalization layers, decreasing initialization scale is equivalent to decreasing the scale of the outputs, since we can write $f(Wx)=a^{-p}f(aWx)$ for any scalar $a$ for any homogenous activation $f$. This in turn is equivalent to increasing the scale of the labels. Therefore, decreasing initialization forces the weights to change more in order to fit the labels, leading to more change in $F$ from its initialization. Conceptually, this aligns with the substantial line of empirical and theoretical evidence that increasing initialization scale or output scaling transitions training between the lazy and feature learning regimes [6, 32, 33, 34].
This relationship suggests that small initialization can be broadly applied to increase the change in F, and the value of the UC-NFC. However, this may not be ideal; for example, if the activation function is differentiable at 00, small initialization leads to a network which is approximately linear. This may lead to low expressivity unless the learning dynamics can increase the weight magnitude.
@@ 5.2 Speed Limited Optimization
Figure 4: The effect of SLO on C/UC neural feature correlations and feature learning on the chain monomial task. In the first two rows, we plot the uncentered and centered NFA for the first layer weight matrix as a function of initialization, (A) with standard training and (B) with SLO. We consider a two hidden layer network with ReLU activations, where we set $C_{0}=500$, and $C_{1}=C_{2}=0.002$. The third column shows the ratio of the unnormalized C-NFC to the UC-NFC: ${\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)\cdot{\text{tr}\left(W^{\top}WW^{\top}KW\right)}^{-1}}$. The fourth column shows the training loss. In the third row, we plot the NFM and AGOP from a trained network with (C) standard training and (D) with Speed Limited Optimization with fixed initialization scale of $1.0$.
We can instead design an intervention which can increase feature learning without the need to decrease the initialization scale. We do so by fixing the learning speed layerwise to constant values, which causes the C-NFC to dominate the UC-NFC dynamics. For weights at layer $\ell$ and learning rate $\eta>0$, we introduce Speed Limited Optimization (SLO), which is characterized by the following update rule,
@@
where the hyperparameter $C_{\ell}\geq 0$ controls the amount of learning in layer $\ell$. We expect this rule to increase the strength of the UC-NFC in layers where $C_{\ell}$ is large relative to $C_{m}$ for $m\neq\ell$, as $W^{(\ell)}$ will be forced to change significantly from initialization. As a result, $\left\|W_{0}^{(\ell)}\right\|\left\|W_{t}^{(\ell)}\right\|^{-1}\rightarrow 0$ for large $t$, causing the centered and uncentered NFC to coincide for this layer.
We demonstrate the effects of SLO on the chain-monomial task (Equation (2)). We found that fixing the learning speed to be high in the first layer and low in the remaining layers causes the ratio of the unnormalized C-NFC to the UC-NFC to become close to $1$ across initialization scales (Figure 4). The same result holds for the SVHN dataset (see Appendix L). We note that this intervention can be applied to target underperforming layers and improve generalization in deeper networks (see Appendix J for details).
We observe that both the C-NFC and the UC-NFC become close to $1$ after training with SLO, independent of initialization scale (Appendix H). Further, the quality of the features learned, measured by the similarity of $F$ and $\bar{G}$ to the true EGOP, significantly improve and are more similar to each other with SLO, even with large initialization. In contrast, in standard training the UC-NFA fails to develop with large initialization, as $F$ resembles identity (no feature learning), while $\bar{G}$ only slightly captures the relevant features.
Speed Limited Optimization is a step toward the design of optimizers that improve generalization and training times through maximizing feature learning.
@@ 6 Discussion
Centered NFC through training
Our theoretical analysis in this work shows that the PTK feature covariance at initialization has a relatively simple structure in terms of the weights and the data. To predict the NFC later in training, we will likely need to account for the change in this matrix. One should be able to predict this development at short times by taking advantage of the fact that eigenvectors of the Hessian change slowly during training [35]. An alternative approach would be to use a quadratic model for neural network dynamics [36, 37] which can capture dynamics of the empirical NTK (and therefore the PTK).
@@
We provide an analytic prediction for the C-NFC at initialization for one hidden layer networks with quadratic activation functions, and provide first-order predictions for two hidden layer networks with quadratic and ReLU activations. A followup direction would be to use random matrix analysis to more accurately predict the NFC for deep architectures with general activations. A promising approach in this direction is to approximate non-linear activations with a linear expansion and independent correction terms [38, 39, 40, 41]. Alternatively, we may hope to establish a general spectral condition on $\mathcal{K}$ that facilitates high correlation between $\dot{W}^{\top}\dot{W}$ and $\dot{W}^{\top}K\dot{W}$ at early training times.
@@
There are prior works that implement differential learning rates across layers [42, 43], though these differ from differential learning speeds as in SLO, in which the L2-norm of the gradients are fixed, and not the scaling of the gradients. Our intervention is more similar to the LARS [44] and LAMB [45] optimizers that fix learning speeds to decrease training time in ResNet and BERT architectures. However, these optimizers fix learning speeds to the norm of the weight matrices, while our intervention sets learning speeds to free hyperparameters in order to increase feature learning, as measured by the NFC. We also note as an additional explanation for the success of SLO in our experiments, the PTK feature covariance may change slowly (or just in a single direction) with this optimizer, because the learning rate is small for later layers, allowing the first layer weights to align with the PTK.
@@
We demonstrate in this work that the dataset (by our construction in Section 4) and optimizer (with SLO in Section 5) play a significant role in the strength of feature learning. Important future work would be to understand other settings where significant empirical differences exist between optimization choices. In particular, analyzing the NFC may clarify the role of gradient batch size and adaptive gradient methods in generalization [46].
Appendix A Glossary
  1. 1.
    Expected gradient outer product (EGOP). Defined with respect to a target function $f^{*}$ and an input data distribution $\mu$ over data in $d$ dimensions.
    $\operatorname{EGOP}(f^{*},\mu)\equiv\mathbb{E}_{x\sim\mu}\left[\nabla_{x}f^{*}(x)\nabla_{x}f^{*}(x)^{\top}\right]\in\mathbb{R}^{d\times d}.$
  2. 2.
    Average gradient outer product (AGOP). Defined with respect to a predictor $f$ and a set of inputs $\{x^{(1)},\ldots,x^{(n)}\}$.
    $\mathrm{AGOP}(f,\{x^{(\alpha)}\}_{\alpha})\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)})}{\partial x}\frac{\partial f(x^{(\alpha)})}{\partial x}^{\top}\in\mathbb{R}^{d\times d}.$
    In the context of a deep neural network, we write the AGOP $\bar{G}_{\ell}$ with respect to the inputs of a given layer $\ell$ as
    $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)})}{\partial h_{\ell}}\frac{\partial f(x^{(\alpha)})}{\partial h_{\ell}}^{\top}.$
  3. 3.
    Neural feature matrix (NFM). Given a neural network $f$ with a weight matrix $W^{(\ell)}$ at layer $\ell$, the NFM $F_{\ell}$ is defined as,
    $F_{\ell}\equiv(W^{(\ell)})^{\top}W^{(\ell)}.$
  4. 4.
    Neural feature correlation (NFC). Defined with respect to a layer of a neural network, this is the correlation $\rho\left(F_{\ell},\bar{G}_{\ell}\right).$
  5. 5.
    Neural feature ansatz (NFA). This refers to the statement that the NFC will have correlation approximately equal to $1$.
  6. 6.
    Pre-activation tangent kernel (PTK). This is the kernel function corresponding, defined with respect to a neural network and a layer $\ell$, that evaluates two inputs $x,z$ to $\frac{\partial f(x)}{\partial h_{\ell}}\cdot\frac{\partial f(z)}{\partial h_{\ell}}\in\mathbb{R}$.
  7. 7.
    PTK feature covariance. For a given layer of a neural network and a dataset $X$, the PTK feature covariance $K^{(\ell)}$ is defined as $K^{(\ell)}\equiv\frac{\partial f(X)}{\partial h_{\ell}}^{\top}\frac{\partial f(X)}{\partial h_{\ell}}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$.
  8. 8.
    Centered NFC (C-NFC). Consider a layer $W$ of a neural network with PTK feature covariance $K$ with respect to that layer. Let $\bar{W}\equiv W-W_{0}$, where $W_{0}$ are weights at initialization. Then, the C-NFC is equivalent to the correlation $\rho\left(\bar{W}^{\top}\bar{W},\bar{W}^{\top}K\bar{W}\right)$.
  9. 9.
    Uncentered NFC (UC-NFC). This quantity is identical to the NFC defined above.
Appendix B Connection of the PTK to the empirical NTK
We note the following connection between the PTK entries and the empirical NTK (ENTK). In particular, the PTK is a significant component of the ENTK.
Proposition 4 (Pre-activation to neural tangent identity).
Consider a depth-$L$ neural network $f(x)$ with inputs $X_{\ell}\in\mathbb{R}^{n\times k_{\ell}}$ to weight matrices $W^{(\ell)}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$ for $\ell\in[L]$. Consider the empirical NTK where gradients are taken only with respect to $W^{(\ell)}$ evaluated between two points $x$ and $z$ denoted by $\hat{\Theta}_{\ell}(x_{0},z_{0})=\langle\frac{\partial f(x_{0})}{\partial W^{(\ell)}},\frac{\partial f(z_{0})}{\partial W^{(\ell)}}\rangle$. Then for all $\ell$, $\hat{\Theta}_{\ell}$ satisfies,
@@ Appendix B Connection of the PTK to the empirical NTK
Appendix C Additional centerings of the NFC
Double-centered NFC
One may additionally center the PTK feature map to understand the co-evolution of the PTK feature covariance and the weight matrices. We consider such a centering that we refer to as the double-centered NFC, where we measure $\rho\left((\bar{W}^{(\ell)})^{\top}\bar{W}^{(\ell)},(\bar{W}^{(\ell)})^{\top}\bar{K}^{(\ell)}\bar{W}^{(\ell)}\right)$, where $\bar{K}^{(\ell)}=\left(\frac{\partial f(X)}{\partial h_{\ell}}-\frac{\partial f_{0}(X)}{\partial h_{\ell}}\right)^{\top}\left(\frac{\partial f(X)}{\partial h_{\ell}}-\frac{\partial f_{0}(X)}{\partial h_{\ell}}\right)$, and $f_{0}$ is the neural network at initialization.
However, the double-centered NFC term corresponds to higher-order dynamics that do not significantly contribute the centered and uncentered NFC (Figure 5) when initialization is large or for early periods of training. Note however this term becomes relevant over longer periods of training.
Figure 5: Ratio of the unnormalized double-centered NFC to the centered NFC throughout neural network training. In particular, we plot $\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}\bar{K}\bar{W}\right)\cdot{\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)}^{-1}$ throughout training for both layers of a two-hidden layer MLP with ReLU activations.
@@ Appendix C Additional centerings of the NFC
One may also center just the PTK feature map, while substituting the initial weights for $W$ to isolate how the PTK feature covariance aligns to the weight matrices. To measure this alignment, we consider the PTK-centered NFC, which is defined as the correlation $\rho\left((W_{0}^{(\ell)})^{\top}W_{0}^{(\ell)},(W_{0}^{(\ell)})^{\top}\bar{K}^{(\ell)}W_{0}^{(\ell)}\right)~{}$, where $W_{0}^{(\ell)}$ is the initial weight matrix at layer $\ell$.
@@
Figure 6: PTK-centered NFC throughout training for both layers of a two-hidden layer MLP with ReLU activations on Gaussian data with two different spectra.
Appendix D Extending our theoretical predictions to depth and general activations
Precise predictions of the C-NFC become more complicated with additional depth and general activation functions. However, we note that the deep C-NFC will remain sensitive to a first-order approximation in which $K$ is replaced by its expectation. We demonstrate that this term qualitatively captures the behavior of the C-NFC for 2 hidden layer architectures with quadratic and, to a lesser extent, ReLU activation functions in Figure 7. In this experiment, we sample Gaussian data with mean 00 and covariance with a random eigenbasis. We parameterize the eigenvalue decay of the covariance matrix by a parameter $\alpha$, called the data decay rate, so that the eigenvalues have values $\lambda_{k}=\frac{1}{1+k^{\alpha}}$. As $\alpha$ approaches 00 or $\infty$ the data covariance approaches a projector matrix.
In this experiment, we see that the data covariance spectrum will also parameterize the eigenvalue decay of $\mathbb{E}\left[K\right]$, allowing us to vary how close the expected PTK matrix (and its dual, the PTK feature covariance) is from a projector, where the NFA holds exactly. We see that for intermediate values of $\alpha$, both the observed and the predicted derivatives of the C-NFC decreases in value.